OpenStudy (bacon):

A jar of cookie is the is 630g when 1/2 full(including the jars weight) and 855g when 7/9 full. what is the mass of the empty jar? how do i solve it?

OpenStudy (anonymous):

\[630=1/2 \] \[855=7/9\] So u can find the ration of that by subtiact the ratio of 7/9 with 1/2 and 855 with 630, that's is \[850-630x=5/18\] Solve that You will get 225g for each 5/18 each1/18 is equal 45g so, multiple by 18, so that's how much a jar can hold that's equal 810

OpenStudy (anonymous):

let jars weight be x and total volume of jar be y now equation y1/2+x=630 7/9y+x=855 sove the two equation by substitution method